OpenStudy (anonymous):

Simplify: (4a^5/3)^3/2

jimthompson5910 (jim_thompson5910):

\[\Large \left(4a^{\frac{5}{3}}\right)^{\frac{3}{2}}\] \[\Large 4^{\frac{3}{2}}a^{\left(\frac{5}{3}*\frac{3}{2}\right)}\] \[\Large 4^{\frac{3}{2}}a^{\frac{5*3}{3*2}}\] \[\Large 4^{\frac{3}{2}}a^{\frac{15}{6}}\] \[\Large 4^{\frac{3}{2}}a^{\frac{5}{2}}\] \[\Large \sqrt{4^3}a^{\frac{5}{2}}\] \[\Large \sqrt{64}a^{\frac{5}{2}}\] \[\Large 8a^{\frac{5}{2}}\] \[\Large 8\sqrt{a^5}\] \[\Large 8\sqrt{a^4*a}\] \[\Large 8\sqrt{a^4}*\sqrt{a}\] \[\Large 8a^2\sqrt{a}\] So \[\Large \left(4a^{\frac{5}{3}}\right)^{\frac{3}{2}}=8a^2\sqrt{a}\]

OpenStudy (anonymous):

Thank you so much!